A certain mobile company offers a monthly plan of Php 799 and an additional charge of Php 7.50 per minute when calling to other networks.How much is the monthly bill if the costumer made a 5 minute call to other networks?
A.What is the independent and dependent variable?
B.What is the slope,y-intercept and the linear inequalities?''.
C.What is the value of y?

Answer by CPhill(2138)

Let's break down this problem.
A. Independent and Dependent Variables
Independent Variable: The number of minutes spent calling other networks (let's call it 'x').
Dependent Variable: The total monthly bill (let's call it 'y').
B. Slope, Y-intercept, and Linear Equation
Fixed Monthly Fee: Php 799
Cost per Minute: Php 7.50
Linear Equation (Slope-Intercept Form):
y = mx + b
Where:
m = slope (cost per minute)
b = y-intercept (fixed monthly fee)
Therefore, the equation is: y = 7.50x + 799
Slope: 7.50 (This represents the cost per minute of calling other networks)
Y-intercept: 799 (This represents the fixed monthly fee)
C. Value of Y (Total Monthly Bill)
The customer made a 5-minute call to other networks (x = 5).
Substitute x = 5 into the equation:
y = 7.50(5) + 799
y = 37.50 + 799
y = 836.50
Answer: The total monthly bill is Php 836.50.